# Course Allocation with Credits via Stable Matching

### v1

@@ Abstract
In the Course Allocation problem, there are a set of students and a set of courses at a given university. University courses may have different numbers of credits, typically related to different numbers of learning hours, and there may be other constraints such as courses running concurrently. Our goal is to allocate the students to the courses such that the resulting matching is stable, which means that no student and course(s) have an incentive to break away from the matching and become assigned to one another. We study several definitions of stability and for each we give a mixture of polynomial-time algorithms and hardness results for problems involving verifying the stability of a matching, finding a stable matching or determining that none exists, and finding a maximum size stable matching. We also study variants of the problem with master lists of students, and lower quotas on the number of students allocated to a course, establishing additional complexity results in these settings.
@@ 1.1 Background and motivation
Every year, undergraduate students must select a list of courses to enrol on for the next academic year. Their selections are constrained by several factors, mainly that university courses have a limited capacity and that universities enforce a limit on the academic load of students. This load is usually expressed in terms of the total number of course credits, with longer courses that have more associated learning hours typically having more credits than shorter ones.
[…]
@@
In some cases, all courses within a faculty have the same preferences over students elte_regulations_2025 . For example, all courses may give priority to students with higher grades over those with lower grades. We show that, when such a uniform ranking of students exists, known as a master list, it is always possible to find a stable matching in polynomial time for all four stability criteria. However, stable matchings can have different sizes, and we show that the problem of finding a stable matching of maximum size is NP-hard, except in the case of pair stability. Finally, we consider the case of courses having lower quotas biro_college_2010 , a minimum number of students that a course needs in order to run. In general we find that problems involving lower quotas are NP-hard.
@@ 1.2 Related work
The Course Allocation problem has been extensively studied in the past, with papers in this area studying a range of different models of the problem. One of the best known mechanisms for addressing this problem is the Approximate Competitive Equilibrium from Equal Incomes mechanism (A-CEEI) budish_combinatorial_2011 , in which students rank possible schedules (each schedule is a subset of courses which does not violate any constraint) and each is given an approximately equal budget to bid on courses. The outcome is approximately Pareto efficient and satisfies some other fairness criteria. It was later refined, resulting in the Course Match mechanism budish_course_2017 . Due to the difficulty in expressing an ordering across all possible schedules, later work has modified Course Match to use machine learning to better estimate the true preferences of students soumalias_machine_2024 .
@@
Alternative approaches to the Course Allocation problem involve students ranking courses individually, instead of ranking bundles, but ensuring that the final allocation respects constraints such as the maximum number of credits a student can take or disallowing excluded combinations such as courses that run concurrently biswas_2023_algorithmic . Other types of constraints have also been considered, for example where courses have prerequisites, corequisites or similar constraints that force students, if they desire to enrol in a given course, to also enrol in a different one cechlarova_pareto_2018 .
[…]
@@
Stable matching problems were introduced by Gale and Shapley gale_college_1962 . They considered the Stable Marriage problem (sm), a problem in which agents are divided into men and women and must be paired together. They show that a stable matching, a matching in which no man-woman pair wants to break with their partners and be together instead, always exists and can be found in polynomial time. The algorithm that they presented is known as the Deferred Acceptance mechanism (DA). The Hospitals/Residents problem (hr) is a variation of sm in which resident doctors apply to hospitals, and hospitals may have more than one resident. As in the case of sm, a stable matching always exists and can be found in polynomial time using DA gale_college_1962 . However, a variant known as Hospitals/Residents with Couples (hrc), in which some residents form a couple and submit a joint preference list, is NP-hard ng_1988_complexity ; ron_1990_np , even if each member of every couple has the same preference list as their partner mcdermid_keeping_2010 . This variant is known as Hospitals/Residents with Sizes, because each couple can be treated as a resident of size $2$ mcdermid_keeping_2010 . In the case of Course Allocation, if every course has the same number of credits, DA can be used to find a stable matching. However, as in the case of couples in hrc, having courses with different numbers of credits means that DA generally does not output a stable matching.
[…]
@@
In many matching markets, including some Course Allocation settings elte_regulations_2025 , agents in one side are ordered according to a master list, such that the preferences of the agents on the other side are consistent with the master list. The presence of a master list guarantees the existence of a stable matching in problems such as Stable Roommates abraham_stable_2007 and makes some NP-hard problems solvable in polynomial time abraham_stable_2007 . However, this is not always the case. Generally, problems such as finding a maximum size stable matching are NP-hard even in the presence of master lists, if there are ties present abraham_stable_2007 ; irving_stable_2008 .
@@
In some applications, courses have lower quotas, representing a minimum number of students that need to be enroled in the course in order for the course to run cechlarova_pareto_2017 , but this extension has not received a great deal of attention in the literature on Course Allocation. However, there are many results for problems such as Hospitals/Residents (sometimes in the equivalent setting of university admissions). Some papers study the case where every hospital must satisfy its lower quota; this case is known as the no closures model. If hospitals have the choice of satisfying the lower quota or being closed and not admitting any residents, the model is the closures model. It is possible to determine in polynomial time whether there exists a stable matching in the hr with lower quotas and no closures model, because in every stable matching each hospital has the same number of residents gale_1985_some ; roth_1984_evolution ; roth_1986_allocation , so one simply needs to check whether in a stable matching all hospitals satisfy their lower quotas. On the other hand, if there are ties in the preference lists of the participants, hr with lower quotas and no closures is NP-complete, because determining the existence of a complete stable matching is NP-complete iwama_2002_stable ; manlove_2002_hard , so setting lower quotas to be equal to upper quotas proves the result for hr with lower quotas and no closures. On the other hand, for hr with lower quotas and closures, the problem of deciding whether a stable matching exists is NP-complete. biro_college_2009 ; biro_college_2010 .
@@
Our model of the Course Allocation can be placed in the context of the wider literature on matching problems with preferences. For example, if courses are assumed to not have preferences over students, then we obtain a model similar to some previously existing Course Allocation models biswas_2023_algorithmic ; budish_course_2017 , or to Capacitated House Allocation manlove_algorithmics_2013 . If both students and courses have preferences, but students can be matched with at most one course, we obtain the Hospitals/Residents problem gale_college_1962 ; mcdermid_keeping_2010 . If on the other hand courses can be matched with at most one student, every course has one or two credits, and neither students nor courses have preferences, we obtain the Matching with Couples model biro_matching_2014 . In this model finding a maximum size matching is NP-hard biro_matching_2014 .
@@
Aziz et al. aziz_stability_2018 analyse a model similar to ours in the context of allocating refugee families to host localities. Some of their stability definitions are similar to ours; we will elaborate on this point later. Another model of interest is due to Hoyer and Stroh-Maraun in the context of the School Choice problem hoyer_stability_2020 . In their model some students may need more resources than others, which is similar to the fact that, in our model, some courses have a higher number of credits than others. They present a (not necessarily polynomial-time) algorithm that finds a stable matching if one exists. Another paper studying the same problem uses a variation of the Top-Trading Cycles algorithm to output a Pareto optimal matching stroh-maraun_weighted_2024 .
@@ 1.3 Our contributions
Finding (-find) NP-h, $\not\exists$ [T3] NP-h, $\not\exists$ [T4] Open P, $\exists$ [T2]
@@
Lower quotas, master list, P, $\not\exists$ [O6] NP-h, $\not\exists$ [O4] NP-h, $\not\exists$ [O4] NP-h, $\not\exists$ [O4]
@@
Table 1: Summary of our contributions. The rows represent the different problems of interest, while the columns refer to the different definitions of stability. The symbol $\exists$ indicates that a matching of this type always exists, while $\not\exists$ indicates that a matching of this type does not always exist. The results included in this paper are referred to by a letter, indicating whether the result is an observation (O), proposition (P) or theorem (T), followed by the observation, proposition or theorem number.
@@ 1.4 Structure of the paper
The remainder of this paper is structured as follows. Section 2 defines the problem model formally. Section 3 presents results on verifying the stability of a given matching, while Section 4 contains results related to finding a stable matching (or reporting that none exists). Section 5 covers results on finding a maximum size stable matching, and Section 6 summarises the results and presents open problems.
@@ 2 Model
A special case of ca called Course Allocation with Master List (ca-ml) occurs when there is a uniform ranking $\succ_{ML}$ of all students, known as a master list, where $s_{i}\succ_{ML}s_{j}$ if $s_{i}$ is placed higher than $s_{j}$ in $ML$. In the presence of a master list, for all $c_{j}\in C$, $c_{j}$ respects the master list in the sense that, for any two students $s_{i},s_{k}\in P(c_{j})$, $s_{i}\succ_{c_{j}}s_{k}$ if and only if $s_{i}\succ_{ML}s_{k}$. Throughout the paper, we assume that both individual preference lists and master lists are strict.
We assume that each course $c_{j}\in C$ has an associated number of credits, denoted by $O(c_{j})$. We extend this notation to a subset $D\subseteq C$ of courses, i.e., $O(D)=\sum_{c_{j}\in D}O(c_{j})$ is the sum of the number of credits of all courses in $D$. Also course $c_{j}$ has an upper quota (or capacity), denoted by $q^{+}(c_{j})$, which gives the maximum number of students that can be assigned to $c_{j}$. Each student $s_{i}$ has a maximum number of credits, denoted $T(s_{i})$, that she can be assigned from the courses she is allocated.
@@
Although in this work we are mostly concerned with a setting where students take any combination of acceptable courses that does not exceed their credit limit, in real life there are other factors that affect the feasibility of a matching. For example, two courses may run at the same time, students may have credit limits per quarter or semester, or at most one out of three courses from a give group may be selected, because their content is very similar. In general, such constraints can be modelled by defining the feasible subsets of courses for each student. An important special case of such subsets are downward-feasible subsets, as defined by Utture et al. utture_student_2019 which we now define here.
Definition 2 (utture_student_2019 ).
@@
For each $s_{i}\in S$, a set of downward-feasible constraints is a subset $F_{i}$ of the power set of $P(s_{i})$ such that, if $D\in F_{i}$, then for any $D^{\prime}\subseteq D$, $D^{\prime}\in F_{i}$. A matching $M$ is feasible if, for every $s_{i}$, $C_{M}(s_{i})\in F_{i}$. Given a matching $M$ and a student $s_{i}$, we say that a course $c_{j}$ is feasible for $s_{i}$ if $(s_{i},c_{j})$ is acceptable and the matching $M\cup(s_{i},c_{j})$ is feasible.
@@
That is, if it is feasible for a student to be matched with a set of courses, then it is feasible to be matched with a subset of these courses. From the definition it is also clear that credit limits of students are a type of downward-feasible constraints, but we will treat credit limits separately. This is because credit limits are more central to the ca problem; almost all universities and ca models have credit limits on students, while other downward-feasible constraints do not always apply. Every hardness result in this paper assumes that there are no downward-feasible constraints, and so they still apply if these are present.
@@
Throughout this paper we will assume that we can check in polynomial time whether a given matching is feasible, and similarly whether a et of feasible constraints is downward-feasible. In general downward-feasible constraints can model many rules that can be represented efficiently (e.g., a student must take at most $60$ credits per semester, or no two courses scheduled at the same time can be assigned, etc.)
@@
We define the size of a matching to be the total number of credits taken by all students, or alternatively, the number of students enroled in courses multiplied by their number of credits. That is, the size of a matching $M$ is $\sum_{s_{i}\in S}\sum_{c_{j}\in C_{M}(s_{i})}O(c_{j})=\sum_{c_{j}\in C}O(c_{j})\cdot|S_{M}(c_{j})|$. We say that a matching $M$ is course-complete if every course is filled to its capacity, that is, for all $c_{j}\in C$, $|S_{M}(c_{j})|=q^{+}(c_{j})$. A matching is student-complete if every student takes as many credits as possible, that is, for all $s_{i}\in S$, $O(C_{M}(s_{i}))=T(s_{i})$. The following observation is then immediate.
@@
Because courses have different numbers of credits, there are several possible definitions of stability. The most natural definition is the following, which is analogous to the standard definition of stability in the Stable Marriage or Hospitals/Residents. It is equivalent to the definitions of stability in Hospitals/Residents with Sizes mcdermid_keeping_2010 and in School Choice papers by Hoyer and Stroh-Maraun hoyer_stability_2020 and Stroh-Maraun stroh-maraun_weighted_2024 . It is also similar to the definition of stability due to Aziz et al. aziz_stability_2018 .
@@
    $c_{j}$ has fewer students assigned to it than its upper quota $q^{+}(c_{j})$ or $c_{j}$ prefers $s_{i}$ over some $s_{k}$ assigned to $c_{j}$. That is, $|S_{M}(c_{j})|<q^{+}(c_{j})$ or $s_{i}\succ_{c_{j}}s_{k}$ for some $s_{k}\in S_{M}(c_{j})$.
@@
    There exists a subset $D_{i}\subseteq C_{M}(s_{i})$ (possibly empty) such that $s_{i}$ prefers $c_{j}$ to every $c_{k}\in D_{i}$, and $O(C_{M}(s_{i}))-O(D_{i})+O(c_{j})\leq T(s_{i})$. That is, $s_{i}$ has enough credits remaining for $c_{j}$, possibly after removing a subset of courses that $s_{i}$ prefers less than $c_{j}$.
@@
    If there are downward-feasible constraints, the matching $(M\cup\{(s_{i},c_{j})\})\setminus\{(s_{i},c_{k}):c_{k}\in D_{i}\}$ is feasible. That is, removing $D_{i}$ and adding $c_{j}$ to $C_{M}(s_{i})$ does not result in an infeasible matching.
If $M$ has no blocking pair then it is pair-stable. Course Allocation with pair stability is denoted ca-p.
An example of a matching with a blocking pair is given in Figure 1. However, we can see that in this example $s_{1}$ would have fewer credits if she was matched with $c_{1}$ instead of $c_{3}$. In such a situation, it may be argued that such a student is not better off, since students generally need to have a curriculum that is as complete as possible. That motivates our next definition of stability. This definition is similar to the definition of stability by demand of Aziz et al. aziz_stability_2018 .
@@
Figure 1: Example of a matching in a Course Allocation instance, highlighted in green, and a blocking pair, highlighted in red.

### v2

# José Rodríguez , David Manlove

@@ Abstract
In the course allocation problem, there are a set of students and a set of courses at a given university. University courses may have different numbers of credits, typically related to different numbers of learning hours, and there may be other constraints such as courses running concurrently. Students may have upper bounds on the number of credits they may be assigned, and courses may have upper quotas on the number of students that can be enrolled. Our goal is to allocate the students to the courses, taking these constraints into account, such that the resulting matching is stable, which means that no student and course(s) have an incentive to break away from the matching and become assigned to one another. We study several definitions of stability and for each we give a mixture of polynomial-time algorithms and hardness results for problems involving verifying the stability of a matching, finding a stable matching or determining that none exists, and finding a maximum size stable matching. We also study variants of the problem with master lists of students, and lower quotas on the number of students allocated to a course, establishing additional complexity results in these settings.
Keywords: Course Allocation problem; credits; master list; lower quotas; stable matching
@@ 1.1 Background and motivation
Every year, undergraduate students must select a list of courses to enrol on for the next academic session. Their selections are typically constrained by several factors, mainly that university courses have a limited capacity and that universities enforce a limit on the academic load of students. This load is usually expressed in terms of the total number of course credits, with longer courses that have more associated learning hours typically having more credits than shorter ones.
These constraints, along with others such as prerequisites or excluded combinations (e.g., due to concurrent timetabling), have led to a rich variety of papers on methods for finding an optimal (with respect to some given criteria) allocation of students to courses Budish (2011); Budish et al. (2017); Budish and Cantillon (2012); Shriya and Ranjan (2026); Sönmez and Ünver (2010). In general, these models assume that students have preferences over bundles of courses, but that courses are indifferent to the students they are allocated. However, this is not the case in many universities, with some giving a higher priority to senior students, or students with better grades Diebold and Bichler (2017); Eötvös Loránd University (Kornbluth and Kushnir, 2024, page 19, footnote 17). An allocation that does not respect these priorities involving students, which can be modelled via course preferences, could lead to justified envy, where a student $s$ who is not already allocated to a course $c$ would rather become allocated to $c$ than remain with their existing assignment, and $s$ has a higher priority at $c$ than some other student already assigned to $c$, or wastefulness, where $s$ would rather be allocated to $c$ than remain with their current assignment, and $c$ has a vacancy. Such a scenario could lead to student $s$ making a complaint, or even persuading the coordinator of course $c$ to change the allocation. This second possibility could lead to further such rearrangements, resulting in the matching unravelling Roth (1984). Hence, we are interested in finding stable matchings, which, informally, are allocations in which there is no justified envy or wastefulness.
Student preferences over bundles of courses can allow constraints such as corequisites and scheduling conflicts to be modelled easily. However, it can be challenging to elicit preferences from students over bundles of courses Budish et al. (2017); Soumalias et al. (2024) and the number of possible bundles is in general exponential in the number of courses. Therefore, in many models Atef Yekta and Day (2020); Biswas et al. (2023); Cechlárová et al. (2018); Diebold and Bichler (2017); Kornbluth and Kushnir (2024); Utture et al. (2019), students have preferences over individual courses, and this is the assumption that we make in this paper. Due to courses having different numbers of credits, stability definitions should be adapted accordingly, and we consider four possible definitions of stability here. All definitions are based on the absence of a blocking pair/coalition, which involves a student and either a single course or a set of courses. In the first definition, pair stability, a student may block a matching (i.e., form a blocking pair) with a single course only, as long as no constraints are violated by the switch. In the second one, pair-size stability, a student may also block with a single course, as long as she does not reduce her total number of credits by making this switch. In the last two definitions, coalition stability and first-coalition stability, each student may block with a coalition of courses, as long as she does not reduce her total number of credits. The difference between them is that, in coalition stability, the student has to prefer every course in the coalition over every course that she wants to drop from the matching, while in first-coalition stability, the student only needs to prefer her favourite course from the coalition over her favourite course among those she wants to drop. For both of these coalition-based definitions, each course in the blocking coalition must have an incentive to take on the student in the blocking coalition. For all stability definitions, we study how to verify the stability of a matching and how to find a stable matching, if one exists, and obtain both hardness results and polynomial-time algorithms.
@@
In some cases, all courses within a faculty have the same preferences over students Eötvös Loránd University . For example, all courses may give priority to students with higher grades over those with lower grades. We show that, when such a uniform ranking of students exists, known as a master list, it is always possible to find a stable matching in polynomial time for all four stability criteria. However, stable matchings can have different sizes, and we show that the problem of finding a stable matching of maximum size is NP-hard, except in the case of pair stability. Finally, we consider the case of courses having lower quotas Biró et al. (2010), a minimum number of students that a course needs in order to run. In general we find that problems involving lower quotas are NP-hard.
@@ 1.2 Related work
The course allocation problem has been extensively studied in the past, with papers in this area studying a range of different models of the problem. One of the best known mechanisms for addressing this problem is the Approximate Competitive Equilibrium from Equal Incomes mechanism (A-CEEI) Budish (2011), in which students rank possible schedules (each schedule is a subset of courses which does not violate any constraint) and each is given an approximately equal budget to bid on courses. The outcome is approximately Pareto efficient and satisfies some other fairness criteria. It was later refined, resulting in the Course Match mechanism Budish et al. (2017). Due to the difficulty that students may have with expressing an ordering across all possible schedules, later work has modified Course Match to use machine learning to better estimate the true preferences of students Soumalias et al. (2024).
@@
Alternative approaches to the course allocation problem involve students ranking courses individually, instead of ranking bundles, but ensuring that the final allocation respects constraints such as the maximum number of credits a student can take or disallowing excluded combinations, such as courses that run concurrently Biswas et al. (2023). Other types of constraints have also been considered, for example where courses have prerequisites, corequisites or similar constraints that force students, if they desire to enrol in a given course, to also enrol in a different one Cechlárová et al. (2018).
Many papers on course allocation consider courses to be objects that do not have preferences or priorities over the students that are allocated to them. However, for many university courses around the world, certain groups of students are given a higher priority than others Diebold and Bichler (2017); Eötvös Loránd University (Kornbluth and Kushnir, 2024, page 19, footnote 17). This is also the case for the School of Law at the University of Glasgow Graves and Wierzbicki (2025), whose course allocation model inspired this work. In particular, when courses have preferences, it is critical that the allocation of students to courses is a stable matching, because stability is a crucial criterion for the success of a matching mechanism in a bipartite matching market with two-sided preferences Roth (1984, 2008).
@@
Stable matching problems were introduced by Gale and Shapley Gale and Shapley (1962). They considered the stable marriage problem (sm), a problem in which agents are divided into men and women and must be paired together. They showed that a stable matching, a matching in which no man-woman pair wants to break with their partners and be together instead, always exists and can be found in polynomial time. The algorithm that they presented is known as the Deferred Acceptance mechanism (DA). The hospitals/residents problem (hr) is a variation of sm in which resident doctors apply to hospitals, and hospitals may be assigned more than one resident. As in the case of sm, a stable matching always exists and can be found in polynomial time using DA Gale and Shapley (1962). However, a variant known as hospitals/residents with couples (hrc), in which some residents form a couple and submit a joint preference list, is NP-hard Ng and Hirschberg (1988); Ronn (1990), even if each member of every couple has the same preference list as their partner McDermid and Manlove (2010). This variant is a special case of hospitals/residents with sizes (hrs), a variant of hr where each resident occupies multiple posts at a hospital, because each couple can be treated as a resident of size $2$ McDermid and Manlove (2010). In the case of course allocation, if every course has the same number of credits, DA can be used to find a stable matching. However, as in the case of couples in hrc, having courses with different numbers of credits means that DA is not guaranteed to output a stable matching.
There has been some previous work on course allocation that considers stability. For example, Kornbluth and Kushnir Kornbluth and Kushnir (2024) introduced a competitive equilibrium-based mechanism where the outcome is approximately stable. However, in their model all courses are assumed to have the same academic load, or number of credits, which is not always the case Biswas et al. (2023); Budish et al. (2017); University of Glasgow . Utture et al. Utture et al. (2019) do consider a setting where courses have different numbers of credits and other constraints may be present, such as excluding combinations of courses that run in the same time slot, but the algorithmic framework they present does not always output a stable matching. Diebold and Bichler Diebold and Bichler (2017) also study course allocation with stability, but in their model students can be matched to at most one course. Hence, this work is the first one that studies course allocation with stability in a model where courses have different numbers of credits, there are excluded combinations of courses, and where certain other constraints may be present.
@@
In many matching markets, including some course allocation settings Eötvös Loránd University , agents in one side are ordered according to a master list, such that the preferences of the agents on the other side are consistent with the master list. The presence of a master list guarantees the existence of a stable matching in problems such as stable roommates Abraham et al. (2008) (a non-bipartite generalisation of sm) and makes some NP-hard problems solvable in polynomial time Abraham et al. (2008). However, this is not always the case. Generally, problems such as finding a maximum size stable matching are NP-hard even in the presence of master lists, if there are ties present Abraham et al. (2008); Irving et al. (2008).
@@
In some applications, courses have lower quotas, representing a minimum number of students that need to be enrolled in the course in order for the course to run Cechlárová and Fleiner (2017), but this extension has not received a great deal of attention in the literature on course allocation. However, there are many results for problems such as hr (sometimes in the equivalent setting of university admissions). Some papers study the case where every hospital must satisfy its lower quota; this case is known as the no closures model. The variant where hospitals have the choice of satisfying the lower quota or being closed and not admitting any residents is known as the closures model. It is possible to determine in polynomial time whether there exists a stable matching in the hr with lower quotas and no closures model, because in every stable matching each hospital has the same number of residents Gale and Sotomayor (1985); Roth (1984, 1986), so one simply needs to check whether in one stable matching all hospitals satisfy their lower quotas. On the other hand, if there are ties in the preference lists of the participants, giving the hospitals/residents problem with ties (hrt), then in the presence of lower quotas and in the no closures model, the problem of deciding whether a stable matching exists is NP-complete. This follows because determining the existence of a complete stable matching is NP-complete Iwama et al. (1999); Manlove et al. (2002) even in the 1-1 restriction of hrt known as smti (sm with ties and incomplete lists), and a complete stable matching can be enforced by setting every lower quota to be $1$. However, for hr with lower quotas and strictly ordered preferences in the closures model, the problem of deciding whether a stable matching exists is NP-complete. Biró et al. (2009, 2010).
@@
Our model of course allocation can be placed in the context of the wider literature on matching problems with preferences Manlove (2013). For example, if courses are assumed to not have preferences over students, then we obtain a model similar to some previously existing course allocation models Biswas et al. (2023); Budish et al. (2017), or to capacitated house allocation Manlove (2013). If both students and courses have preferences, but students can be matched with at most one course, we obtain the hr problem Gale and Shapley (1962); McDermid and Manlove (2010). If on the other hand courses can be matched with at most one student, every course has one or two credits, and neither students nor courses have preferences, we obtain the matching with couples model Biró and McDermid (2014). In this model finding a maximum size matching is NP-hard Biró and McDermid (2014).
In the literature on matching under preferences, there have been several papers considering problems where agents have different sizes. For example, Balasundaram et al. Balasundaram et al. (2025) studied hrs; we will return to this paper in Section 1.3. Delacrétaz Delacrétaz (2019) considered a version of matching with different sizes in which agents have size $1$ or $2$, and presented different relaxations of stability that ensure the existence of a stable matching. One of them, size-stability, is equivalent to pair-size stability.
@@
Aziz et al. Aziz et al. (2018) analysed a model similar to ours in the context of allocating refugee families to host localities, which is similar to our model as refugee families can be regarded as agents with different sizes. Some of their stability definitions are similar to ours; we will elaborate on this point later, in Section 2. Delacrétaz et al. Delacrétaz et al. (2023) presented algorithms for finding Pareto optimal, individually rational or interference-free (a concept related to envy-freeness) matchings in the same context. Another model of interest is due to Hoyer and Stroh-Maraun in the context of the school choice problem Hoyer and Stroh-Maraun (2020). In their model some students may need more resources than others, which is similar to the fact that, in our model, some courses have a higher number of credits than others. They presented a (not necessarily polynomial-time) algorithm that finds a stable matching if one exists. Another paper studying the same problem uses a variation of the Top-Trading Cycles algorithm to output a Pareto optimal matching Stroh-Maraun (2024).
@@ 1.3 Our contributions
Finding (-find) NP-h, $\nexists$ [T3] NP-h, $\nexists$ [T4] Open P, $\exists$ [T2]
@@
Lower quotas, master list, P, $\nexists$ [O7] NP-h, $\nexists$ [O4] NP-h, $\nexists$ [O4] NP-h, $\nexists$ [O4]
@@
Table 1: Summary of our contributions. The rows represent the different problems of interest, while the columns refer to the different definitions of stability. The symbol $\exists$ indicates that a matching of this type always exists, while $\nexists$ indicates that a matching of this type need not always exist. The complexity results included in this paper are denoted by ‘P’ for polynomial-time solvability, ‘NP-h’ for NP-hardness and ‘co-NP-c’ for co-NP-completeness. They are also followed by a letter, indicating whether the result is an observation (O), proposition (P) or theorem (T), followed by the observation, proposition or theorem number.
Independently of our results, Balasundaram et al. Balasundaram et al. (2025) recently considered hrs under two definitions of stability: firstly stability, which is equivalent to pair stability in our model, and secondly occupancy-stability, which is equivalent to pair-size stability. They showed that an occupancy-stable matching is guaranteed to exist, and can be found in polynomial time, which is equivalent to Theorem 2, and proved that it is NP-hard to find a maximum size occupancy-stable matching, a similar result to Theorem 8, although in our theorem students and courses’ preferences are derived from a master list, while in their result there is no master list. Balasundaram et al. Balasundaram et al. (2025) did not consider constraints such as excluded combinations.
On the other hand, they provide results that do not appear in our work, such as an approximation algorithm for finding a maximum size occupancy-stable (pair-size-stable) matching with performance guarantee strictly better than $3$, and a tighter reduction proving that it is NP-complete to determine whether an instance of hrs has a stable (pair-stable) matching. Finally, they also consider the problem with master lists. Although their definition of a master list is not equivalent to ours, they also show that a stable matching always exists, for any definition of stability, as we do in Proposition 2. Hence, their paper Balasundaram et al. (2025) and the preliminary version of our paper Rodríguez and Manlove (2025), published independently of each other and in short succession, have partially overlapping results, but each considers slightly different problems and contains several original results not present in the other.
@@ 1.4 Structure of the paper
The remainder of this paper is structured as follows. Section 2 defines the problem model formally. Section 3 presents results on verifying the stability of a given matching, while Section 4 contains results related to finding a stable matching (or reporting that none exists). Section 5 covers results on finding a maximum size stable matching (both in the absence of, and in the presence of, lower quotas), and Section 6 summarises the results and presents open problems.
@@ 2 Model
A special case of ca called course allocation with master list (ca-ml) occurs when there is a uniform ranking $\succ_{MLS}$ of all students, known as a master list, where $s_{i}\succ_{MLS}s_{j}$ if $s_{i}$ is placed higher than $s_{j}$ in the uniform ranking. In the presence of a master list, for all $c_{j}\in C$, $c_{j}$ respects the master list in the sense that, for any two students $s_{i},s_{k}\in P(c_{j})$, $s_{i}\succ_{c_{j}}s_{k}$ if and only if $s_{i}\succ_{MLS}s_{k}$. Throughout the paper, we assume that both individual preference lists and master lists are strict.
We assume that each course $c_{j}\in C$ has an associated number of credits, denoted by $O(c_{j})$. We extend this notation to a subset $D\subseteq C$ of courses, i.e., $O(D)=\sum_{c_{j}\in D}O(c_{j})$ is the sum of the number of credits of all courses in $D$. Also course $c_{j}$ has an upper quota, denoted by $q^{+}(c_{j})$, which gives the maximum number of students that can be assigned to $c_{j}$. Each student $s_{i}$ has a maximum number of credits, denoted $T(s_{i})$, that she can be assigned from the courses she is allocated.
@@
Given a matching $M$, if $C_{M}(s_{i})=\emptyset$ for a student $s_{i}\in S$, we will say that $s_{i}$ is unmatched in $M$. Similarly if $S_{M}(c_{j})=\emptyset$ for a course $c_{j}\in C$, we will say that $c_{j}$ is unmatched in $M$. If $|S_{M}(c_{j})|<q^{+}(c_{j})$ for a course $c_{j}\in C$, we will say that $c_{j}$ is undersubscribed in $M$.
@@
Although in this work we are mostly concerned with a setting where students take any combination of acceptable courses that does not exceed their credit limit, in real life there are other factors that affect the feasibility of a matching. For example, two courses may run at the same time (they are a so-called excluded combination), students may have credit limits per quarter or semester, or at most one out of three courses from a given group may be selected, because their content is very similar. In general, such constraints can be modelled by defining the feasible subsets of courses for each student. An important special case of such subsets are downward-feasible subsets, as defined by Utture et al. Utture et al. (2019) which we now define here.
Definition 2 (Utture et al. (2019)).
@@
For each $s_{i}\in S$, a set of feasible constraints is a subset $F_{i}$ of the power set of $P(s_{i})$. $F_{i}$ is downward-feasible if $D\in F_{i}$ and $D^{\prime}\subseteq D$ implies that $D^{\prime}\in F_{i}$. A matching $M$ is feasible if, for every $s_{i}$, $C_{M}(s_{i})\in F_{i}$. Given a matching $M$ and a student $s_{i}$, we say that a course $c_{j}$ is feasible for $s_{i}$ if $c_{j}\in P(s_{i})$ and $C_{M}(s_{i})\cup\{c_{j}\}\in F_{i}$.
@@
That is, downward feasibility means that if it is feasible for a student to be matched with a set of courses, then it is feasible to be matched with a subset of these courses. From the definition it is also clear that credit limits of students are a type of downward-feasible constraints, but we will treat credit limits separately. This is because credit limits are more central to the ca problem; almost all universities and ca models have credit limits on students, while other downward-feasible constraints do not always apply. Every hardness result in this paper assumes that there are no downward-feasible constraints, and so they still apply if these are present.
@@
Throughout this paper we will assume that we can check in polynomial time whether a given matching is feasible, and similarly whether a set of feasible constraints is downward-feasible. In general downward-feasible constraints can model many rules that can be represented efficiently (e.g., a student must take at most $60$ credits per semester, or no two courses scheduled at the same time can be assigned, etc.)
@@
We define the size of a matching to be the total number of credits taken by all students, or alternatively, the number of students enrolled in courses multiplied by their number of credits. That is, the size of a matching $M$ is $\sum_{s_{i}\in S}\sum_{c_{j}\in C_{M}(s_{i})}O(c_{j})=\sum_{c_{j}\in C}O(c_{j})\cdot|S_{M}(c_{j})|$. We say that a matching $M$ is course-complete if every course is filled to its upper quota, that is, for all $c_{j}\in C$, $|S_{M}(c_{j})|=q^{+}(c_{j})$. A matching is student-complete if every student takes as many credits as possible, that is, for all $s_{i}\in S$, $O(C_{M}(s_{i}))=T(s_{i})$. The following observation is then immediate.
@@
Because courses have different numbers of credits, there are several possible definitions of stability. The most natural definition is the following, which is analogous to the standard definition of stability in sm and hr. It is equivalent to the definitions of stability in hrs Balasundaram et al. (2025); McDermid and Manlove (2010) and in school choice papers by Hoyer and Stroh-Maraun Hoyer and Stroh-Maraun (2020) and Stroh-Maraun Stroh-Maraun (2024). It is also similar to the definition of stability due to Aziz et al. Aziz et al. (2018).
@@
    $c_{j}$ is undersubscribed or $c_{j}$ prefers $s_{i}$ over some $s_{k}$ assigned to $c_{j}$. That is, $|S_{M}(c_{j})|<q^{+}(c_{j})$ or $s_{i}\succ_{c_{j}}s_{k}$ for some $s_{k}\in S_{M}(c_{j})$.
@@
    There exists a subset $D_{i}\subseteq C_{M}(s_{i})$ (possibly empty) such that $s_{i}$ prefers $c_{j}$ to every $c_{k}\in D_{i}$, and $O(C_{M}(s_{i}))-O(D_{i})+O(c_{j})\leq T(s_{i})$. That is, $s_{i}$ has enough credits remaining for $c_{j}$, possibly after removing a subset of courses that $s_{i}$ prefers $c_{j}$ to.
@@
    If there are downward-feasible constraints, $(C_{M}(s_{i})\setminus D_{i})\cup\{c_{j}\}\in F_{i}$. That is, removing $D_{i}$ and adding $c_{j}$ to $C_{M}(s_{i})$ does not result in an infeasible matching.
If $M$ has no blocking pair then it is pair-stable. ca with pair stability is denoted ca-p.
An example of a matching with a blocking pair is given in Figure 1. However, we can see that in this example $s_{1}$ would have fewer credits if she was matched with $c_{1}$ instead of $c_{3}$. In such a situation, it may be argued that such a student is not better off, since students generally need to have a curriculum that is as complete as possible. That motivates our next definition of stability. This definition is similar to the definition of stability by demand due to Aziz et al. Aziz et al. (2018), and equivalent to the occupancy-stability definition due to Balasundaram et al. Balasundaram et al. (2025) and the size-stability definition due to Delacrétaz Delacrétaz (2019).
@@
Figure 1: Example of a matching in a ca instance, highlighted in green, and a blocking pair, highlighted in red.